A 500g wooden block can slide down an inclined track uniformly when the inclination is 28 ° from the horizontal. What is the magnitude of the friction force?
1
Expert's answer
2020-06-29T14:05:20-0400
"F_{net}=0=F_f-mg\sin{28}"

"F_f=(0.5)(9.8)\sin{28}=2.3\ N"